OpenStudy (anonymous):

The total farm population (the number of people living on farms ) in the United States was 6,051,000 in 1980. Since then, it has been decreasing at the rate of 118,250 people per year. How many people were living on farms in 1990?

OpenStudy (saifoo.khan):

Make an equation. As the popl is decreasing every year, the slope of your equation will be 118,250

OpenStudy (saifoo.khan):

Haha, yeah i know many people copy my last name.. They all are obsessed with it. :3

OpenStudy (anonymous):

okay got that so: 118,250x+y

OpenStudy (anonymous):

rite?

OpenStudy (anonymous):

wait a minute

OpenStudy (anonymous):

from 1980 to 1990 its ten years so the slope times 10 or is it slope times -10

OpenStudy (saifoo.khan):

y = mx + b y = 118250x + b

OpenStudy (saifoo.khan):

WAIT! it should be -ve.

OpenStudy (anonymous):

so x is 10 rite> or is it -10

OpenStudy (anonymous):

-ve?

OpenStudy (saifoo.khan):

Why do you think it's -10?

OpenStudy (saifoo.khan):

-ve = negative.

OpenStudy (anonymous):

um rong its postive ten well i thought its -10 because it says decreasing

OpenStudy (anonymous):

so what should it be? how can I get the answer

OpenStudy (saifoo.khan):

y = mx + b Where slope wil be: 118,250 And since the popl is decreasing it will be -118,250.

OpenStudy (anonymous):



OpenStudy (saifoo.khan):

Hence: y = -118,250x + b now find the value of b. Insert the given point: (x,y) = (1980, 6051000) y = -118,250x + b 6051000 = -118,250(1980) + b Solve for b.

OpenStudy (saifoo.khan):

PS b should be a huge number.

OpenStudy (saifoo.khan):

Okay let me tell you. b = 240186000

OpenStudy (anonymous):

wait I got a diffrentn answer

OpenStudy (anonymous):

oh my bad

OpenStudy (anonymous):

so the y=mx+b would be y=-118,250x + 240186000

OpenStudy (anonymous):

but we didnt answer the question

OpenStudy (anonymous):

wait don't we substitue 10 for x

OpenStudy (anonymous):

hello?

OpenStudy (saifoo.khan):

Yes right.

OpenStudy (anonymous):

about substtuting 10 for x

OpenStudy (anonymous):

:)

OpenStudy (saifoo.khan):

Haha. Okay sorry. Miss T?

OpenStudy (anonymous):

its okay its the first and last time

OpenStudy (anonymous):

dont call me that either

OpenStudy (saifoo.khan):

Okay unknown person.

OpenStudy (saifoo.khan):

Lol. No Mrs now.

OpenStudy (anonymous):

so back to the problem i subtutte rite?

OpenStudy (saifoo.khan):

Yes 10/

OpenStudy (anonymous):

I know thanks